# ToMAP: Training Opponent-Aware LLM Persuaders with Theory of Mind

### v1

@@ 3.2 A Reinforcement Learning Formulation for Persuader Training
We adopt proximal policy optimization (PPO) [59], a widely-used and stable algorithm, to train the persuader policy $\mathbf{X}$, optimizing it with the standard PPO objective function:
$\max_{\mathbf{X}}\mathbb{E}_{{x\sim\mathcal{D},T\sim\mathbf{X}}}\Big{[}r(T,Q)-\beta\mathbb{D}_{\text{KL}}\big{(}\mathbf{X}(T\mid x)\ \big{|}\big{|}\ \mathbf{X}_{\text{ref}}(T\mid x)\big{)}\Big{]}$ (7)
where $D$ is the dataset, $r$ is the reward function, $\beta$ is a hyperparameter controlling KL penalty and $\mathbf{X}_{\text{ref}}$ is the reference policy of the persuader. Notably, $x=(p,Q,H)$ contains the system prompt, target claim, as well as the conversation history. Overall, the RL schema enables the persuader to explore diverse strategies, learn from conversational signals, and ultimately develop effective persuasion techniques.
3.3 Theory of Mind (ToM) Modules
Theory of Mind (ToM)—the ability to understand others’ mental states, especially opinions and attitudes—is essential for effective persuasion. To equip LLMs with such ability and provide crucial information for their strategic persuasion, we incorporate Counterclaim Predictor and Opponent Attitude predictor. into our persuader agent, as illustrated in Figure 2.
3.3.1 Counterclaim Predictor
Unlike human persuaders, current large language models (LLMs) often struggle to conceptualize a debate topic and its associated claims as a structured, interconnected cognitive graph. This limitation reduces their ability to anticipate an opponent’s beliefs and potential counterarguments. Therefore, we introduce the counterclaim predictor module, a novel component designed to encourage the persuader to proactively anticipate counterarguments that might be held by the persuadee. Based on that, the persuader can craft more resilient and contextually aware persuasive strategies.
$\neg q_{1\ldots k}\sim\mathbf{X}(p,Q,k),$ (8)
3.3.2 Opponent Attitude Predictor
While the counterclaim predictor effectively identifies potential objections, it cannot capture the nuanced shifts in opinion that occur during a conversation. To provide the persuader with such a dynamic understanding of the persuadee’s evolving beliefs, we propose the opponent attitude predictor to predict the persuadee’s opinion on opposing claims generated by Equation 8. By leveraging the attitude predictor, the persuader can gain a more accurate understanding of the persuadee’s thoughts, which are necessary for flexible and targeted persuasion in the subsequent turns.
Implementation. As zero-shot prompting doesn’t yield satisfactory results, we use an external classifier, which consists of a text encoder and a multilayer perceptron (MLP), to predict opponent attitudes. Specifically, we use the pre-trained BGE-M3 encoder [60] to encode the claim and the conversation as vectors separately. The two embedding are then concatenated and passed to a 5-way MLP classifier, which predicts the agreement score $s^{\prime}$:
where $\mathbf{E}$ denotes the encoder, $\mathbf{M}$ denotes a multi-layer perceptron, $H$ is the conversation history and $Q$ is the claim. “||||| |” refers to the concatenation of vectors. With the cross-entropy loss, we train $\mathbf{M}$ to accurately predict the persuadee’s attitudes on counterclaims generated by the counterclaim predictor. Refer to Section A.2 for details regarding the attitude predictor.
3.3.3 Incorporating ToM Information
Information from the ToM modules is included in the persuader’s prompt during training and inference, serving as an augmentation of input $x$ in the PPO formulation (Equation 7):
$x_{\text{aug}}=(x,\neg q_{1\ldots k},s^{\prime}_{\text{pre}}(H,\neg q_{1\ldots k}))=\big{(}p,Q,H,\neg q_{1\ldots k},s^{\prime}_{\text{pre}}(H,\neg q_{1\ldots k})\big{)}.$ (10)
In conclusion, the ToMAP framework trains the persuader with ToM-aware reinforcement learning, enabling it to effectively incorporate the opponent’s mental state information into its persuasive planning, resulting in more diverse and impactful arguments.
4 Experiments
4.1 Experimental Settings
Datasets. During training, we use the Cornell CMV dataset [61], a widely used corpus derived from Reddit’s ChangeMyView forum. It encompasses a broad range of topics, including political views, philosophy, environment, daily habits, and more. We train the persuader on 18.2k claims. In evaluation, we also included claims derived from the Anthropic Persuasion Dataset [20] and the args.me corpus [62] to assess persuaders in out-of-distribution persuasion topics. Notably, all topics collected are controversial issues, so we set the persuadee’s initial stance to be negative in the prompt to ensure $S(\mathbf{Y},H_{0},Q)<0.5$. Please refer to Appendix B for prompts, Appendix C for data statistics and Appendix F for persuadees’ initial attitudes.
Base Models. We train persuader models based on QWen-2.5-3B-Instruct [63], and the persuadee model during training is QWen-2.5-7B-Instruct. During evaluation, we also evaluate persuaders against LLaMa-3.1-8B-Instruct [64] and Phi-4  [65], to assess the generalizability.
Metric. We use agreement shift, the gap between the persuadee’s agreement score before and after the conversation, as the metric indicating persuasiveness. Formally, the agreement shift is calculated as $S(\mathbf{Y},H_{2n},Q)-S(\mathbf{Y},H_{0},Q)$ ($n=3$ in the main experiment), which is then scaled to a 0–100 range and reported as a percentage. During evaluation, we set the temperature to $1.0$ to encourage diverse generation, and report the average score over 3 trials to ensure evaluation stability [^3] The standard deviation across the three trials is at most 0.97. A one-tailed t-test indicates a statistically significant difference between RL and ToMAP with $p<0.01$..
Experimental Settings. We evaluate five baseline LLMs and two finetuned persuaders:
•  Baseline LLMs: QWen-2.5-3B-Instruct (referred as Base), Gemma-3-27B-it [66], LLaMa-3.1-70B-Instruct [64], GPT-4o [67] and Deepseek-R1 [55].
•  RL: the base model trained using reinforcement learning (RL) without incorporating Theory of Mind (ToM) information. Please refer to Appendix D for details regarding the training procedure.
Table 1: ToMAP outperforms baselines in persuasiveness. Numbers represent the agreement shift in the conversation where two agents each take 3 turns. The best results in each row are bolded.
Dataset Persuadee Persuader (Size in B)
Base Gemma-3 Llama-3.1 Deepseek-R1 GPT-4o RL ToMAP
3 27 70 671 N/A 3 3
CMV QWen-7B 12.75 13.52 12.58 16.04 13.36 16.42 23.01
LLaMa-8B 4.97 11.35 2.23 8.51 3.35 15.15 8.82
Phi-4 (14B) 19.44 33.09 19.84 32.31 23.97 12.27 20.93
Anthropic QWen-7B 5.66 10.02 5.27 7.81 7.62 8.92 15.82
LLaMa-8B 7.09 7.54 4.10 8.79 3.90 19.59 12.10
Phi-4 (14B) 23.24 29.91 20.50 30.61 26.56 14.84 22.79
args.me QWen-7B 6.90 12.21 7.09 10.34 9.51 10.77 18.75
LLaMa-8B 4.09 9.69 2.22 7.69 2.50 14.89 10.89
Phi-4 (14B) 21.74 28.07 21.29 31.13 22.06 11.71 24.19
Average N/A 11.76 17.27 10.57 17.02 12.54 13.84 17.48
4.2 Main Results
Results in Table 1 highlight the following findings:
RL is effective in enhancing persuasiveness. As the Base model isn’t trained or designed to be a persuader, its influence on the opponent’s opinion is very limited. Through training on carefully designed reward signals that encourage highly persuasive arguments, RL shows a remarkable relative performance gain of 17.78% compared with the Base model.
ToM modules make stronger persuaders. Equipped with ToM modules that model the opponent’s mental state, ToMAP shows a further relative gain of 26.14% compared to RL. This advantage stems from ToMAP’s ability to explicitly model the opponent’s thoughts and attitudes on the cognition graph. These predictive capabilities allow for a deeper understanding of the context and the opponent, ultimately enhancing their persuasiveness. Notably, while ToMAP model only contains 3B parameters, it outperforms much larger baselines [^4] ToMAP’s advantages over Gemma-3 and Deepseek-R1 are marginal; however, given the significantly larger parameter counts of these models, the results still highlight ToMAP’s impressive persuasiveness., indicating that small language models can also be persuasive through proper training. Additionally, experiments in Appendix A show that the ToM modules excel in modeling opponent thoughts, further validating our approach.
ToMAP generalizes well to diverse scenarios. Among the 9 combinations of persuadee and corpus, ToMAP outperforms RL on 6 combinations, and outperforms Base on 8. In contrast, other persuaders’ performances vary greatly across different persuadees: Deepseek-r1 influences Phi-4’s opinion greatly, and RL performs the best against LLaMa-8B, but their persuasion outcomes against other persuadees are limited. These results underscore the crucial role of the attitude predictor in enabling persuaders to learn more flexible persuasion strategies that persuade opponents based on their thoughts.
4.3 Ablation Studies
4.3.1 RL is Necessary to Incentivize Effective Persuasion
Figure 3: Zero-shot prompting with ToM information doesn’t enhance the performance of the baseline model. This figure compares Base and ToMAP in two settings: without and with ToM information in the user prompt. Titles of subgraphs indicate corpus and persuadee.
As described in Section 3.3, the information from the counterclaim predictor and the attitude predictor is prepended to the user prompt for the persuader, which doesn’t inherently require finetuning the persuader model. This raises a natural question: can the ToM modules enhance the base model’s performance in a zero-shot manner? Our experiments reveal that the answer is negative, indicating that the persuadee’s mental state information can only demonstrate its full effectiveness after RL.
From Figure 3, we can clearly observe that the optimal performance is achieved only in the “ToMAP model with ToM information” setting. Notably, merely adding ToM information to the Base model without proper training yields minimal performance gains and can even be detrimental in some tasks, as unusable information often acts as a distraction. In addition, removing the ToM information from the ToMAP model also degrades its performance in most cases, since ToMAP can no longer access the information it has learnt to use during training. Therefore, both ToM modules and RL training are necessary for the performance gain: ToM modules provide valuable assessments of the opponent, while RL enables the persuader to exploit this information and develop a better strategy accordingly.
4.3.2 The Attitude Predictor is Crucial for Persuasiveness
Since the attitude predictor is a key component in ToMAP, we evaluate the impact of ablating this component in this subsection. Specifically, we compare ToMAP with several variants that incorporate different Theory of Mind (ToM) information: ToMAP(w/o att) is trained without the opponent attitude predictor and only uses the counterclaim predictor [^5] ToMAP(w/o att) is trained separately, unlike the other variants which have same parameters as ToMAP.; ToMAP-rand uses random opponent attitudes; ToMAP-max sets all opponent attitudes to be completely agree; and ToMAP-gt uses the ground truth opponent attitudes obtained by querying the persuadee.
Table 2: ToMAP outperforms baseline settings with missing or inaccurate attitude information, and approaches the performance of the ground-truth attitude setting. The persuadee here is QWen-2.5-7B-Instruct.
Setting CMV Anthropic args.me
Base 12.75 5.66 6.90
ToMAP(w/o att) 15.57 9.55 10.45
ToMAP-rand 21.63 14.66 17.17
ToMAP-max 21.69 16.14 16.55
ToMAP-gt 23.57 17.05 16.65
ToMAP 23.01 15.82 18.75
From Table 2, we can find that ToMAP outperforms all the settings with missing or inaccurate opponent attitudes, demonstrating the critical role of our “Encoder+MLP” based predictor in effectively modeling opponent attitude. For the other settings, ToMAP(w/o att), which fails to anticipate opponent attitudes, performs the worst; ToMAP-rand and ToMAP-max, which inject potentially untruthful attitudes, also underperform ToMAP, suggesting that inaccurate predictions may hinder finding the optimal persuasion strategy. Notably, the minor gap between ToMAP and the upper bound ToMAP-gt further confirms the predictor’s capability to estimate the opponent’s attitudes reliably.
5 Analysis
5.1 Understanding ToMAP through Training Process
(a) Persuasion Reward
(b) Repetition Penalty [^7] Repetition penalty is part of the reward design, aiming to measure lexical overlap between multiple persuader turns. Refer to Appendix D for details.
(c) Thought Length (tokens)
(d) Argument Length (tokens)
Figure 4: Plots for key metrics during RL training. We apply Time Weighted Exponential Moving Average (EMA) with $\alpha=0.3$ to smooth the plots (the semi-transparent lines are data before smoothing).
We show the trend of important metrics during training for RL, ToMAP (w/o att), and ToMAP in Figure 4, which reveal the following insights:
Figure 4(a) demonstrates that all settings exhibit remarkable gains during the RL process, and the persuasion rewards of all persuader policies converge within 200 training steps, indicating that the RL training process effectively unlocks the full potential of each model. Notably, ToMAP achieves a significantly higher reward than the baselines from the 90th step, aligning with the benchmarking results reported in Table 1.
As shown in Footnote 7, the standard RL setting suffers from increased repetition toward the later stages of training. In contrast, ToM-enhanced variants maintain high argumentative diversity throughout the entire training process. This highlights the benefit of proactively modeling the opponent’s possible counterclaims, which contributes to generating more varied and engaging arguments, instead of focusing solely on the central claim and repeating already-expressed arguments.
As shown in Figures 4(c) and 4(d), while all settings maintain similar argument lengths, ToMAP shows a clear trend of increasing thought length compared to other models. This is consistent with the “long thought” phenomenon in [55], which is an indicator of complex reasoning, implying that planning the persuasion strategy with ToM information is a reasoning-intensive task. It also reinforces the conclusion in Section 4.3.1, that effective analysis over the opponent’s mental state is a capability that emerges only through reinforcement learning.
5.2 ToMAP Achieves Stable Persuasion Gains Over Turns
77footnotetext: Legend only shown once for clarity.
From Figure 5, we observe that while RL initially outperforms the baseline in the first three conversational turns, it fails to continuously increase the agreement score in subsequent rounds, with performance plateauing and even declining. In contrast, ToMAP demonstrates a steady and substantial improvement across turns, culminating in a notable 10.64 extra agreement shift compared to RL, and 11.86 compared with Base by the 10th turn. The consistent persuasion gain suggests that ToMAP’s design enables dynamic adaptation and opponent-aware strategic refinement as conversations progress. The effective modeling of the opponent’s thoughts and attitudes makes ToMAP especially suitable for long and complex persuasive exchanges.
(a) CMV99footnotemark: 9
(b) Anthropic
(c) args.me
Figure 5: ToMAP shows steady gains in long conversations. The persuadee is QWen-2.5-7B-Instruct.
Figure 6: Strategy usage comparison between the base model and ToMAP. ToMAP adopts a more reasoned, audience-aware strategy than the base model.
5.3 Persuasion Techniques of ToMAP
Humans naturally employ a wide variety of persuasion strategies that help them convince the opponent and achieve the persuasion goal. Although ToMAP was not explicitly programmed with any specific persuasion techniques, it was trained through RL to explore effective strategies autonomously. Motivated by this, we experiment to compare the persuasion techniques that are frequently utilized by Base and ToMAP in practice.
Following prior work [68, 69], we define a taxonomy of 9 persuasion strategies (see Appendix E) and use GPT-4o to annotate their usage in Base and ToMAP. As shown in Figure 6, ToMAP more frequently employs “Common Ground”, “Preemptive Rebuttal” and “Concession”, indicating a more opponent-aware approach. It relies less on strategies that resort to sentiments, like “Rhetoric” and “Social Appeals”, favoring logical reasoning. Slightly more frequent use of “Evidence” and reduced “Authority Appeal” further support that ToMAP adopts a more logical, strategic and opponent-aware argumentative style, benefiting from the ToM information about the opponent’s thoughts. An illustrative example of ToMAP’s argument is provided in Appendix E.
6 Conclusion
In this paper, we introduce Theory of Mind Augmented Persuader (ToMAP), a novel framework designed to imbue language models with enhanced persuasive capabilities by explicitly modeling the persuadee’s mental state. ToMAP incorporates a counterclaim predictor, enabling the anticipation of potential objections, and an opponent attitude predictor, facilitating an understanding of the persuadee’s agreement levels on related claims. Our carefully constructed reinforcement learning schema empowers the persuader to effectively utilize these ToM-driven insights for generating more diverse and impactful arguments.
Extensive experiments showed that ToMAP consistently outperformed several baselines, including much larger state-of-the-art models, across diverse scenarios. Further analysis revealed that ToMAP fosters more complex reasoning and reduces repetition, leading to more diverse and effective arguments. Notably, ToMAP shows consistent persuasion gains in long conversations where its audience awareness enables more logical and opponent-aware strategy design. In conclusion, this work marks a promising attempt towards developing more human-like and effective AI persuaders by integrating theory of mind, the crucial principle in human cognition and persuasion process.
Acknowledgment
We extend special thanks to Zirui Cheng for the helpful discussions.
Limitation
In early experiments, we find the persuader often resorts to abnormal “tricks” to get a high reward. This includes prematurely claiming agreement, fabricating evidence (such as nonexistent scientific papers), and presenting broad outlines without elaboration (to fit in the context length). While prompt engineering and reward shaping were employed to mitigate these issues, they represent heuristic solutions rather than fundamental safeguards. Establishing a more resilient framework for training and evaluating persuasive agents against such "tricks" remains a significant area for future research.
Furthermore, the training of ToMAP was conducted using a single persuadee model, QWen-2.5-7B-Instruct. Although our results indicate reasonable generalization to other persuadees (Table 1), incorporating a more diverse set of persuadee models during training could potentially enhance the robustness and adaptability of the learned persuasive strategies.
Finally, we only explored moderate-scale LLMs (3B) as persuaders due to resource constraints. Investigating ToMAP framework’s effect to larger LLMs, with their enhanced knowledge and reasoning capabilities, presents an exciting avenue for future investigation.
Ethical Statement
We acknowledge that persuasive models, if deployed in human-facing systems without robust safeguards, carry the significant risk of being misused to manipulate opinions or propagate misinformation. A recent field study on Reddit CMV [^10] https://www.reddit.com/r/changemyview/comments/1k8b2hj/meta_unauthorized_experiment_on_cmv_involving (the same source used in the main paper), while controversial, emphasizes such ethical risks.
To rigorously ensure safety, our development process incorporates multiple layers of protection. This includes expert scrutiny to proactively identify and mitigate unintended persuasive effects of the model. Furthermore, we are committed to a careful and controlled release of our models, prioritizing safety at each stage. Especially, we will release a comprehensive usage guideline together with the model, designed to actively discourage malicious applications. Finally, We also believe transparency is the key to avoid misuse of persuader models. Specifically, users should explicitly know when AI is persuading them and have the option to opt out. We strongly discourage “turing-test” style persuasion where the identity of the AI is unknown, since this could result in irreversible opinion shift.
Appendix A Details about Theory of Mind Modules
This section provides implementation details about the two theory of mind modules used in the main paper, and conducts experiments to demonstrate their effectiveness.
A.1 Counterclaim Predictor
Table 3: Cosine similarity between QWen-2.5-3B-Instruct and different persuadees’ claims. Claims generated by different LLMs are semantically similar.
Persuadee Similarity (%)
QWen-7B 73.72
LLaMa-8B 73.82
Phi-4 (14B) 74.72
@@ A.1 Counterclaim Predictor
Specifically, we compare the persuader model QWen-2.5-3B-Instruct and three persuadee models, as detailed in Section 4.1. We prompt different LLMs to generate 3 claims supporting each topic in the CMV test set. We then use the BGE-M3 text encoder to convert these claims into vector representations and compute the cosine similarity between them to measure their semantic similarity. To fairly compare sets of claims between two models, we apply the Hungarian Algorithm [70] to identify the optimal one-to-one mapping that maximizes overall similarity. This allows us to assess how well the persuader’s generated counterclaims align with the perspectives that persuadee models would articulate themselves.
From Table 3, we can find that different models’ claims for supporting a claim are semantically similar, with similarity scores larger than 73%. Furthermore, the case study in Table 4 confirms that different models, like QWen-2.5-3B-Instruct and Phi-4, focus on similar aspects of the topic. These results suggest that prompting the persuader to propose counterclaims is effective in modeling the opponent’s mental state.
Table 4: An example of counterclaim predictor (based on QWen-2.5-3B-Instruct) and Phi-4 (the real persuadee)’s claims on the same topic. From the comparison, we can find the two LLMs focus on similar points.
@@
A.2 Attitude Predictor
@@
Table 5: The MLP predictor outperforms LLM prompting in modeling opponent attitudes. The metrics are exact match (EM) and mean square error (MSE) [^12] We adopt the discrete 5-way division of attitudes in Section 3.1 where the range is $\{0,1,2,3,4\}$.. Higher EM and lower MSE indicates better performance.
@@
Hyper-parameters. As mentioned in Section 3.3, the task of the attitude predictor can be formulated as $s^{\prime}_{\text{pre}}(H,Q)=\mathbf{M}(\mathbf{E}(H)\|\mathbf{E}(Q))$, where $\mathbf{E}$ is the trainable MLP. In our implementation, the MLP has 3 intermediate layers whose sizes are 1024, 256, and 64, respectively, connected by the ReLU activation function. We train the MLP with a learning rate of $45e-$ 4 for 20 epochs, record validation loss after each epoch, and keep the checkpoint with the best performance on the validation set. We finalize the above hyperparameters through grid search.
@@
Performance. We analyse the performance of our attitude predictor and compare it with zero-shot prompting LLMs [^13] When prompting the LLM, we provide the LLM with the conversation history and target claim, and then ask it to choose from the five attitudes after chain-of-thought reasoning.. From Footnote 12, we can observe that our predictor reaches decent accuracy and has a relatively small mean square error (MSE) compared with the baselines, highlighting its effectiveness in modeling the persuadee’s evolving attitudes. Additionally, the MLP has only $\sim$4M parameters, and the backbone encoder is also lightweight with $\sim$0.5B parameters, ensuring inference-time computational efficiency.
Appendix B Prompts
In this section, we show all the prompts used in the experiments.
B.1 System Prompts
B.2 User Prompt in the Debate
B.3 Theory of Mind Information
For ToMAP, an additional “Theory of Mind” information is prepended to the user prompt. We show an example for the ToM information in ToMAP setting below. The opponent claims and opponent agreement scores are obtained from the counterclaim predictor and attitude predictor, as mentioned in Section 3.3 and Section A.2.
Appendix C Details about Persuasion Datasets
Table 6: Statistic of datasets and links to sources.
Dataset License # Train # Val
CMV MIT 18.2k 450
Anthropic CC BY-NC-SA 4.0 N/A 75
args.me CC BY 4.0 N/A 400
The CMV dataset comprises posts and comments from the Reddit subreddit r/ChangeMyView, where users present their opinions and invite others to challenge them. The args.me corpus is a large collection of over 380k arguments extracted from four online debate portals: Debatewise, IDebate.org, Debatepedia, and Debate.org. The Anthropic Persuasion Dataset is a collection of claims paired with arguments, both human-written and generated by language models, designed to measure persuasiveness. Table 6 shows the license and amount of data used in this paper.
Since our evaluation criteria involve obtaining the persuadee’s attitude on two contradictory statements (refer to Section 3.1), and the datasets only provide one single claim (or topic), we use GPT-4o to generate claims for both sides. Below is the prompt used to generate the claims:
Appendix D Details about Persuader Training
Table 7: PPO Training Configuration for ToMAP.
Hyperparameter Value Hyperparameter Value
Actor learning rate $1\times 10^{-6}$ Critic learning rate $2\times 10^{-6}$
Warmup ratio $0.2$ Rollout temperature $1.0$
KL Coefficient ($\beta$) $0.001$ Train batch size $128$
PPO mini batch size $64$ PPO micro batch size $32$
Training steps $200$ Max input length $3000$
Max response length $1000$ Max argument length $200$
Number of counterclaims $3$ Turns of persuasion $3$
Our code is based on the verl framework [71] and TinyZero [72], which use Apache-2.0 licenses. RL, ToMAP-lite, and ToMAP are trained in the same configuration, as specified in Table 7. We use 4 NVIDIA RTX A6000 GPUs for training the persuader model, and 1 additional A6000 GPU for deploying the persuadee model with vLLM. Each training step (with 3 turns from the persuader and 3 turns from the persuadee) takes approximately $400$ seconds in our environment. We acknowledge that our hyperparameter choices may not be optimal, as conducting a comprehensive grid search would be computationally prohibitive. However, we ensure a fair comparison across different reinforcement learning settings and find that all settings yield reasonable and relatively stable results.
Appendix E Details about Persuasion Strategies

### v2

@@ C.1 Reward Design
C.2 Training Details
Our code is based on the verl framework Sheng et al. (2024) and TinyZero Pan et al. (2025), which use Apache-2.0 licenses. RL, and ToMAP are trained in the same configuration, as specified in Table˜5. We use 4 NVIDIA RTX A6000 GPUs for training the persuader model, and 1 additional A6000 GPU for deploying the persuadee model with vLLM. Each training step (with 3 turns from the persuader and 3 turns from the persuadee) takes approximately $400$ seconds in our environment. We acknowledge that our hyperparameter choices may not be optimal, as conducting a comprehensive grid search would be computationally prohibitive. However, we ensure a fair comparison across different reinforcement learning settings and find that all settings yield reasonable and relatively stable results. The SFT baseline uses 5k conversations and is trained for 1 epoch with a learning rate of $1\times 10^{-5}$.
C.3 The Persuasion Pipeline
We provide a Python-style pseudocode to illustrate the process of persuasive conversation. The parameter is_ToMAP indicates whether ToM modules are included or not. The pseudocode makes some simplification about attitude calculation, refer to Section˜3.1 for details.
1def Persuasion(persuader, persuadee, n_turns, claim, is_ToMAP=False, attitude_predictor=None):
2 conv_history = []
3 initial_attitude = JudgeAttitude(claim, persuadee)
4
5 for turn in range(n_turns):
6 prompt = FormatPrompt(claim, conv_history)
7 if is_ToMAP:
8 counter_claims = GetCounterClaim(claim)
9 predicted_attitudes = GetOpponentAttitude(conv_history, counter_claims, attitude_predictor)
10 prompt += counter_claims + predicted_attitudes
11 persuader_argument = Inference(persuader, prompt)
12 conv_history.append(persuader_argument)
13
14 prompt = FormatPrompt(claim, conv_history)
15 persuadee_argument = Inference(persuadee, prompt)
16 conv_history.append(persuadee_argument)
17
18 new_attitude = JudgeAttitude(claim, conv_history, persuadee)
19 return conv_history, new_attitude - initial_attitude
Appendix D Details about Theory of Mind Modules
D.1 Counterclaim Predictor
@@
Table 6: An example of counterclaim predictor (based on QWen2.5-3B-Instruct) and Phi-4 (the real persuadee)’s claims on the same topic.
@@
Table 7: Cosine similarity between QWen2.5-3B-Instruct and different persuadees’ claims. Claims generated by different LLMs are semantically similar.
Persuadee Similarity (%)
QWen-7B 73.72
LLaMa-8B 73.82
Phi-4 (14B) 74.72
Specifically, we compare the persuader model QWen2.5-3B-Instruct and three persuadee models, as detailed in Section˜4.1. We prompt different LLMs to generate 3 claims supporting each topic in the CMV test set. We then use the BGE-M3 text encoder to convert these claims into vector representations and compute the cosine similarity between them to measure their semantic similarity. To fairly compare sets of claims between two models, we apply the Hungarian Algorithm Kuhn (1955) to identify the optimal one-to-one mapping that maximizes overall similarity. This allows us to assess how well the persuader’s generated counterclaims align with the perspectives that persuadee models would articulate themselves.
From Table˜7, we can find that different models’ claims for supporting a claim are semantically similar, with similarity scores larger than 73%. Furthermore, the case study in Table˜6 confirms that different models, like QWen2.5-3B-Instruct and Phi-4, focus on similar aspects of the topic. These results suggest that prompting the persuader to propose counterclaims is effective in modeling the opponent’s mental state.
D.2 Attitude Predictor
@@
Hyper-parameters. As mentioned in Section˜3.3, the task of the attitude predictor can be formulated as $s^{\prime}_{\text{pre}}(H,Q)=\mathbf{M}(\mathbf{E}(H)\|\mathbf{E}(Q))$, where $\mathbf{E}$ is the trainable MLP. In our implementation, the MLP has 3 intermediate layers whose sizes are 1024, 256, and 64, respectively, connected by the ReLU activation function. We train the MLP with a learning rate of 5e45e-4 for 20 epochs, record validation loss after each epoch, and keep the checkpoint with the best performance on the validation set. We finalize the above hyperparameters through grid search.
Table 8: The MLP predictor outperforms LLM prompting in modeling opponent attitudes. The metrics are exact match (EM) and mean square error (MSE) [^10] We adopt the discrete 5-way division of attitudes in Section˜3.1 where the range is $\{0,1,2,3,4\}$.. Higher EM and lower MSE indicates better performance.
@@
Performance. We analyse the performance of our attitude predictor and compare it with zero-shot prompting LLMs [^11] When prompting the LLM, we also provide the LLM with the conversation history and target claim, and then ask it to choose from the five attitudes with chain-of-thought.. From Footnote˜10, we can observe that our predictor reaches decent accuracy and has a relatively small mean square error (MSE) compared with the baselines, highlighting its effectiveness in modeling the persuadee’s evolving attitudes. Additionally, the MLP has only $\sim$4M parameters, and the backbone encoder is also lightweight with $\sim$0.5B parameters, ensuring inference-time computational efficiency.
Appendix E Details about Different Persuadees
The prompt design in Appendix˜A makes sure the persuadee disagrees with the target claims initially, leaving ample room for persuasion. Specifically, the persuadees’ average initial agreement scores on the three benchmarks are shown in Table˜9, which confirms our prompt effectively set the original attitudes.
Table 9: Initial scores for each persuadee across datasets.
Persuadee CMV Anthropic args.me
QWen-7B 25.29 33.00 31.66
LLaMa-8B 20.45 18.00 23.66
Phi-4 19.92 24.33 20.53
In Section˜4, we observe that the performance of all persuaders varies across persuadee models. We believe the difference stems from the unique conversation styles of thinking paradigms each model learns during pre-training and post-training. In this section, we present a typical case where ToMAP attempts to convince three target models on the same claim but with different outcomes. Conversations are shortened for readability.
Qwen2.5 as persuadee, score=0.375 CLAIM: Raising the minimum wage to \$15’h will increase unemployment and hurt small businesses. Persuader: Increasing the minimum wage to \$15 places heavy burdens on small businesses, particularly in labor-intensive industries like retail and hospitality. These firms may be forced to reduce employment in order to survive. Persuadee: That is a valid concern, but evidence shows that moderate increases in the minimum wage have not consistently led to job losses. Many businesses adapt through small price adjustments, efficiency improvements, or reduced turnover, which helps maintain stability. Persuader: Nevertheless, low-wage industries are especially fragile. If labor costs rise sharply, their profitability could collapse, resulting in widespread unemployment and business closures. Persuadee: It is possible, but businesses can also invest in technology or adopt new practices that allow them to offset higher wages. Moreover, higher pay creates a more stable and productive workforce, which can reduce training costs and improve competitiveness. Persuader: Without additional policy support, such as targeted subsidies or tax breaks, many small businesses will still struggle to survive. The economic disruptions should not be underestimated. Persuadee: While additional policy support is possible, I agree that such policies must be carefully implemented. Raising the minimum wage abruptly could lead to undesired results.
Phi-4 as persuadee, score=0 CLAIM: Raising the minimum wage to \$15’h will increase unemployment and hurt small businesses. Persuader: Small businesses are the backbone of the economy, and history shows that when the minimum wage rises, job opportunities for low-skilled workers often decline. A sharp increase to \$15 will therefore harm both businesses and employment. Persuadee: That may be true in some cases, but higher wages can also generate significant economic benefits. They lead to greater consumer spending and improve worker morale, which in turn helps small businesses through increased sales and lower turnover. Persuader: These challenges are not temporary fluctuations but systemic shifts. Case studies demonstrate that after wage hikes, unemployment effects can persist for years, especially among the most vulnerable workers. Persuadee: On the other hand, many studies highlight positive long-term outcomes. A living wage can reduce turnover and training expenses, while gradual wage increases and supportive policies can ease the transition for small businesses. Persuader: Current economic conditions, such as inflation and rising costs, make it even more difficult for businesses to absorb higher wages. The likelihood of layoffs is therefore quite high. Persuadee: Even with those pressures, numerous reports show that wage increases often stimulate demand. Balanced policies that combine higher wages with business support can create a more sustainable outcome.
LLaMa3.1 as persuadee, score=-0.25 CLAIM: Raising the minimum wage to \$15’h will increase unemployment and hurt small businesses. Persuader: Raising the minimum wage to \$15 will inevitably harm small businesses. Since these businesses often operate on very thin profit margins, the additional labor costs will likely force them to lay off workers or even close down. Persuadee: I do not fully agree with that conclusion. Many small businesses have shown an ability to adapt to market changes. They may raise prices slightly or improve efficiency to cover the additional costs. Furthermore, several studies suggest that wage increases reduce poverty and reliance on public assistance while stimulating consumer spending. Persuader: While some adaptation is possible, industries such as healthcare or education already operate with high labor costs. Combined with global inflation, these sectors are particularly vulnerable and cannot easily pass higher wages onto customers without suffering severe losses. Persuadee: Yet empirical research, including work from the Federal Reserve, indicates that employment often does not decline after minimum wage increases. In some cases, job opportunities even expand. This evidence challenges the assumption that higher wages necessarily destroy jobs. Persuader: Still, inflation and global market pressures make it extremely difficult for small businesses to adjust smoothly. Under such conditions, layoffs or closures remain a likely outcome. Persuadee: Even so, broader data, such as findings from the Economic Policy Institute, show that higher wages can generate long-term economic growth. From this perspective, the benefits appear to outweigh the risks.
Appendix F Details about Persuasion Strategies